OpenStudy (anonymous):

simplify f(x)=ln((e^x sqroot (x^2+1))/(x-2)^3)

OpenStudy (anonymous):

You can write that as: \[f(x)=\ln(e^x \sqrt{x^2+1})-\ln((x-2)^3)=\ln(e^x)+\ln((x^2+1)^{1/2})-\ln((x-2)^3)\] From here: \[f(x)=x+(1/2)\ln(x^2+1)-3\ln(x-2)\]

OpenStudy (anonymous):

omg do you know how to do f'(x)

OpenStudy (anonymous):

Well, to differentiate that. Go left to right. \[f'(x)=1+(1/2)(2x/(x^2-1))-3/(x-2)\] Since d/dx(ln(x))=1/x

OpenStudy (anonymous):

easy once you simplify the log. may want to write middle term as \[\frac{x}{x^2-1}\]

OpenStudy (anonymous):

mv, how come your latex looks so much nicer than mine?

OpenStudy (anonymous):

Exactly^^ That comes from multiplying the (1/2) by 2x which leaves you with just an x in the numerator. I have no idea D: I'm fairly new here. o.o

OpenStudy (anonymous):

wait the answer is only x

OpenStudy (anonymous):

Nah, its what I have. Satellite was just point out that you could simplify it a little.